For a given location in Ghana,the wind speed at 10m and 50m height above ground has been measured to be4m/s and 6.8m/s respectively. In your own analysis what might be the reason for difference in wind speed
1
Expert's answer
2020-08-26T02:54:46-0400

In wind energy studies, the log law have generally been used to model the vertical profile of wind speed over regions of homogenous, flat terrain. This approach has its origins in boundary layer flow in fluid mechanics and in atmospheric research. It is based on a combination of theoretical and empirical research. 

The increase of wind speed with height in the lowest 100m can be described by this logarithmic expression:


"v=v_{ref}\frac{\ln{(zz_0^{-1})}}{\ln{(z_{ref}z_0^{-1})}}"

where:

  • V = velocity to be calculated at height z
  • Z = height above ground level for velocity v
  • Vref = known velocity at height zref
  • Zref = reference height where vref is known
  • Z0 = roughness length in the current wind direction